OpenStudy (anonymous):

If x varies directly with y and x = 2.5 when y = 10, find x when y = 16. Can someone please help explain this to me?? I don't understand it at all....

OpenStudy (anonymous):

\[x=ky \] solve for k....

OpenStudy (anonymous):

2.5= k(10).......k=4........x=4(16)... so x=4???? Please respond back and tell me if i did this right

OpenStudy (anonymous):

do you get what i mean?

OpenStudy (anonymous):

k =1/4...so u can get x=4..

OpenStudy (anonymous):

Ok you lost me

OpenStudy (anonymous):

you get it???if the statement is varies directly it means x=ky...

OpenStudy (anonymous):

So what I did was right??